# GDP nowcasting with artificial neural networks: How much does long-term memory matter?

### 19 Dec 2023 v2

@@ Abstract
In our study, we apply artificial neural networks (ANNs) to nowcast quarterly GDP growth for the U.S. economy. Using the monthly FRED-MD database, we compare the nowcasting performance of five different ANN architectures: the multilayer perceptron (MLP), the one-dimensional convolutional neural network (1D CNN), the Elman recurrent neural network (RNN), the long short-term memory network (LSTM), and the gated recurrent unit (GRU). The empirical analysis presents the results from two distinctively different evaluation periods. The first (2012:Q1 – 2019:Q4) is characterized by balanced economic growth, while the second (2012:Q1 – 2022:Q4) also includes periods of the COVID-19 recession. According to our results, longer input sequences result in more accurate nowcasts in periods of balanced economic growth. However, this effect ceases above a relatively low threshold value of around six quarters (eighteen months). During periods of economic turbulence (e.g., during the COVID-19 recession), longer input sequences do not help the models’ predictive performance; instead, they seem to weaken their generalization capability. Combined results from the two evaluation periods indicate that architectural features enabling for long-term memory do not result in more accurate nowcasts. On the other hand, the 1D CNN has proved to be a highly suitable model for GDP nowcasting. The network has shown good nowcasting performance among the competitors during the first evaluation period and achieved the overall best accuracy during the second evaluation period. Consequently, first in the literature, we propose the application of the 1D CNN for economic nowcasting.
@@
Journal of Economic Literature (JEL) codes: C45, C53, C51
@@ 1 Introduction
Gross Domestic Product (GDP) is arguably the most important flow-type monetary measure of economic activity over a given period. GDP is often calculated on an annual basis, but in most advanced economies it is also measured on a quarterly frequency. Considering its relevance, the availability and reliability of these quarterly data are of great significance.
@@
In our study, we apply different artificial neural networks (ANNs) to nowcast the quarterly U.S. GDP growth of the current quarter. Using the monthly FRED-MD database, we measure the nowcasting performance of five different ANN architectures relative to a naive constant growth model: The multilayer perceptron (MLP), the one-dimensional convolutional neural network (1D CNN), the Elman recurrent neural network (RNN), the long short-term memory network (LSTM), and the gated recurrent unit (GRU). These architectures are similar in that they all try to meet the above-mentioned challenges of nowcasting. However, the models have a different structure. First, the MLP’s relatively simple architecture is not even designed for sequence modeling. While we can form “training sequences” by hand and add them as input, the MLP does not recognize the temporal ordering in the data. This is not to say that the MLP cannot be used effectively in applications related to time series analysis, but in this regard, the MLP is the exception. Compared to the other ANNs investigated in this paper, the MLP is a simple, time-agnostic architecture, which is not tolerant to shifts (translations) in time ([37], [41]). The 1D CNN and all other RNNs investigated in this paper are specifically designed for sequence modeling. They are capable of learning order dependence in sequence prediction problems. In other words, they are trained with real input sequences. Despite this similarity, even the 1D CNN and the RNNs differ substantially in how they handle the sequential data. [^3] We will see more about these differences in Section 3. From the perspective of the empirical analysis, the question comes down to what architectural features are advantageous in terms of GDP nowcasting.
@@
The first main contribution of this paper is that it proposes the utilization of a neural network architecture that has yet to be applied for economic nowcasting, namely the one-dimensional convolutional neural network (hereinafter, 1D CNN). Our results suggest that the 1D CNN could be a highly suitable architecture for GDP nowcasting as it generates accurate nowcasts for both evaluation periods, i.e., both during a balanced growth period and in times of high economic turbulence. While this ANN architecture has been neglected compared to the MLP and the more advanced gated RNNs, we would argue that it represents a “sweet spot” for this type of analysis in many aspects. In contrast to the MLP, the 1D CNN has an intrinsic architectural constraint that usefully limits the hypothesis space, i.e., the range of mappings the network can learn during the training process. During training, the network learns the coefficients of different finite impulse response filters, so it can only learn a composition of time-invariant (translation-invariant) mappings. Intuitively, the network performs a series of hierarchical pattern recognition tasks where the result is time-invariant to the relative position of those patterns [41]. It is worth noting that the vast majority of the traditional time series approaches (e.g., the dynamic factor model) are also built around this assumption. Namely, the conditional distribution of the target variable only depends on time through the regressors (predictors). From this point of view, it seems especially advantageous to narrow the hypothesis space to this type of mappings. Unlike the MLP, the 1D CNN is trained with input sequences, reflecting the architecture’s temporal “awareness”. Training of the 1D CNN is also helped by the perception of the temporal ordering in input data, ultimately leading to less trainable parameters (weights).
@@
The major novelty of gated RNNs is that they introduce a so-called gating mechanism (practically, element-wise multiplication) that enables the network to develop long-term memory. This feature helps to overcome the vanishing gradient problem of the vanilla RNNs, making them a much more viable option for modeling time series with path-dependent behavior, i.e., with long-term memory ([33], [44]). Financial time series, such as stock prices or even more foreign exchange rates, generally show characteristics of path dependency (e.g., a slowly decaying autocovariance function). Here, it can easily happen that the current value of the exchange rate is affected (supported) by a distant technical level whose effect, however, is no longer reflected in any nearby values. As a result, this old technical level should be remembered by the network in a predictive exercise. Since gated RNNs can learn those long-term dependencies between input and output, they can be highly effective in such predictive exercises. Formation of long-term memory comes at a cost, however: They contain much more trainable weights than an Elman RNN with the same number of neurons in their (hidden) state vector. Consequently, they usually need a large number of training samples to obtain a good level of generalization capability ([37], [41]).
@@
The second main contribution of the paper is that, first in the literature, it presents the results of a comprehensive nowcasting competition between different ANN architectures. The competitor ANN architectures are then trained with different training configurations. These training configurations differ (i) in the length of the training sequences and (ii) in information sets available intra-quarterly. The nowcasting competition presented in this paper has three important aspects and purposes. First, it aims to identify the most suitable ANN architectures for GDP nowcasting. The second research question we try to answer is how important long-term memory is in GDP nowcasting? So, we investigate how the different competitor models’ nowcasting performance changes with different lengths of input sequences. Finally, we also try to measure how the nowcasting accuracy of the competitors changes with consecutive intra-quarterly data releases. So, we distinguish three different nowcasting scenarios depending on how many months of regressor data are available for the models within a given quarter. Accordingly, e.g., the 1-month nowcasting scenario refers to an information set that contains data for the monthly regressors until the first of a given quarter. We believe that the results of the empirical analysis might be indicative for some aspects of the underlying data-generating process as well. If, for example, the data-generating process of GDP growth was highly path-dependent (i.e., non-ergodic), then that would likely speak for the use of those gated RNNs with long input sequences.
@@
We present the results from two evaluation (test) periods during the empirical analysis. While those evaluation periods overlap in time, they are very different in the characteristics of the target series, i.e., GDP growth. The first evaluation period ranges from 2012:Q1 to 2019:Q4, so it ends before the economic consequences of the COVID-19 pandemic would hit the U.S. economy. The results for this period indicate that longer input sequences may lead to more accurate nowcasts for the majority of the competitor models. However, this effect ceases above a relatively low threshold value of around six quarters (eighteen months). While the MLP works best with 8 long sequences in every nowcasting scenario, other competitors yield better accuracy with 18 long input sequences in some nowcasting scenarios. While the 1D CNN produces its best (lowest) relative RMSE and MAE values with 18 long input sequences, the overall best accuracy in the first test period is achieved by the GRU trained with 8 long sequences. Both models beat the benchmark model in terms of nowcasting performance on a 5% significance level. Our second evaluation period is from 2012:Q1 to 2022:Q4, so it also includes those periods of the COVID-19 crisis. Under such circumstances, shorter input sequences lead to more accurate nowcasts for all competitor models. In the second evaluation period, the best accuracy is achieved by the 1D CNN trained with 8 long sequences. We see that architectural complexity clearly hinders the gated RNNs’ nowcasting performance in the second evaluation period. This time, they produce less accurate nowcasts than even the basic Elman RNN. Combined results from the two evaluation periods show that longer input sequences generally do not improve our models’ predictive accuracy. That is especially interesting to see in the case of the LSTM and the GRU: While their architectural design enables them to handle long input sequences, forming a long-term memory, even these networks tend to have better accuracy with shorter sequences. Considering the results from the two test periods as a whole, architectural features supporting the learning of long-term dependencies do not play an important role in GDP nowcasting. Instead, they represent a comprising trade-off between complexity and generalization capability for this specific predictive analysis.
@@ 3 Models and methods
This is not to say however, that ANNs are generally superior modeling frameworks compared to traditional linear models. While ANNs handle the curse of dimensionality more easily, they can perform worse in other areas. For example, they require a much more sophisticated data pre-processing step where we must deal with missing observations as well. One possible solution to fill up missing values is applying univariate ARMA(p,q) forecasts for each single input series. In this case, the lag lengths of the individual ARMA(p,q) models can be selected based on the value of Akaike or Schwarz information criterion ([42], [51]). Although missing data issue for ANNs can be solved in many ways, state space models handle this problem by design.
If we think about the three big challenges of nowcasting mentioned in the introduction, we can say that mixed-frequency data can be treated relatively well in both modeling frameworks. Compared with the DFM, proper construction of our training and test samples with mixed-frequency data may require some additional effort.
More generally, we should also mention that even the smallest ANNs have much more parameters than a typical statistical model. The large number of trainable (free) parameters results in a relatively large hypothesis space. From the bias-variance trade-off follows that with a larger hypothesis space, it is more likely to find a hypothesis which fits the training data very well. This phenomenon is called overfitting, and it happens when a model learns the fine details and noise in the training data rather than the actual systematic relationship between the regressors and the target variable(s). Overfitting can substantially weaken a statistical model’s generalization capability, and ANNs are substantially more prone to overfitting compared to the traditional time series models.
So far, we tried to summarize the most important characteristics of ANNs as one of our nowcasting machines. In the following sections, we investigate those architectures in more detail that were adopted in the empirical analysis.
@@ 3.3 Recurrent neural networks
Due to vanishing gradients, RNNs tend to have a very “short” memory, limiting their usefulness in many applications [37]. Especially in those that require to learn long-term dependencies between inputs and outputs. Even with these considerations, we included one specific variant of the basic RNN in our nowcasting contest: the Elman network or Elman RNN. In an Elman network, each recurrent layer applies the following mapping for each element in the input sequence ([38]):
@@
After the short description of the Elman RNN, we present how to overcome the main limitations (vanishing gradients) associated with the basic RNN by using more advanced units when building the network. Introducing our last two competitor ANN architectures, we shift our focus on more advanced RNN variants that implement a so-called gating mechanism, namely the long short-term memory network and the gated recurrent unit ([44], [33]).
@@
Gated recurrent unit (hereafter, GRU) represents the other gated recurrent network adopted in our nowcasting analysis. Considering its architecture, GRU is very similar to an LSTM but belongs to a newer generation of gated RNNs [33]. The GRU removes the cell state (denoted by $c_{t}$ in the LSTM) and uses its hidden state (denoted by $h_{t}$, similar to the LSTM) even for transferring information. This way, it only has two gates, a reset gate and an update gate. Thereby, GRUs has fewer parameters and tensor operations compared to LSTMs [33]. Consequently, they are a little speedier to train, and a bit less prone to overfitting. Researchers and engineers usually try both to determine which one works better for their specific use case. So we do the same in empirical analysis. Similarly to the 1D CNN, all the RNNs presented above / investigated are trained with sequential data.
@@ 4 Empirical analysis
In this section, we first describe the data used in the empirical analysis then define the concrete nowcasting exercise. After that, we present the results of the empirical analysis.
@@ 4.1 Data and feature selection
The target variable of the empirical analysis is the quarterly GDP growth, that is the percent change (relative to the preceding period) of the US Gross Domestic Product. The series is measured on a quarterly frequency and is available from 1947:Q2 to 2022:Q3. Thus it contains 302 observations. Figure 6 plots the time series for quarterly GDP growth.
Figure 6: The percent change (relative to the preceding period) of the US Gross Domestic Product. Source: Federal Reserve Economic Data (FRED).
When the empirical analysis was closed, the last GDP measurement was released for 2022:Q4. In our empirical analysis, we try to extract the informational content of several monthly macroeconomic indicators (i.e., features) whose intra-quarterly development might be expressive of the current economic growth. The data for the input features comes from the FRED-MD database, the monthly database for Macroeconomic Research of the Federal Reserve Bank of St. Louis, which is described extensively in [54].
FRED-MD is a large macroeconomic database that is designed for the empirical analysis of “big data”. The database is publicly available and updated in real-time on a monthly basis.  [^13] The FRED-MD database is available for download under the following link: https://research.stlouisfed.org/econ/mccracken/fred-databases/. It consists of 134 monthly time series (indicators or features) which are clustered into eight categories:
@@
For the empirical analysis, all monthly indicators (features) are detrended using the corresponding data transformations. This is important because all of our competitor models, except for the MLP, are designed to estimate time-invariant prediction rules, which is preferred for stationary time series. Consequently, when both the target variable and the regressors (features) are stationary, we can generally expect better predictive performance from these models compared to when modeling non-stationary series. Figure 7 below shows the level values of Real personal income (upper panel) and the result of the adequate data transformation (lower panel).
(a) Real personal income (RPI): Original times series with trend.
(b) Real personal income (RPI): Log-differenced, detrended times series.
Figure 7: Real personal income (RPI). Source: FRED-MD.
For the training of the different ANNs, we use all the monthly indicator series included in the FRED-MD database. At the same time, we introduce a bottleneck (encoder) layer, as the first layer in each network, to reduce the dimensionality of the feature space. The resulting low-dimensional embedding (state) vector contains a linear combination of the original variables, so in this respect, it performs a transformation on the data similar to principal component analysis (PCA). In the case of PCA, however, the weights of the resulting linear combination are determined to minimize the reconstruction error. Concretely, the weights used in constructing the $i$-th principal component are the elements of the eigenvector belonging to the $i$-th largest eigenvalue of the covariance matrix. Compared to this, in the solution we propose, the weights of the linear combination are trained in the same integrated training process. Thus they are determined to minimize the prediction error. Considering the relatively small number of available training samples, this type of dimension reduction should significantly help each ANN’s generalization capability. Besides dimension reduction, we also apply regularization to the weight vector for the bottleneck layer ($\mathbf{w}_{b}$). Equation (6) describes the cost function augmented with the an L1 (Lasso) regularization term [47]:
@@
where $L()$ is the loss (criterion) function selected for training, $\lVert\mathbf{w}_{b}\rVert_{1}$ denotes the L1 norm of $\mathbf{w}_{b}$ and $\lambda$ is a hyperparameter which determines how severe the penalty is. [^14] We will specify the loss function later, during the description of the training process. Intuitively, a higher value of $\lambda$ induces a bigger penalty to the L1 norm thus it leads to a more sparse weight vector. As Equation (6) suggests, we do not use regularization in any other layer of of the different networks. This L1 (Lasso) regularization leads to a sparse weight vector in the bottleneck (encoder) layer and ultimately results in feature selection [47]. Figure 8 below shows the histograms of the estimated parameters (weights and biases) for the bottleneck layer ($\mathbf{w}_{b}^{*}$) without regularization (7(a)) and with L1 regularization(7(b)).

### 29 Feb 2024 v3

@@ Abstract
We apply artificial neural networks (ANNs) to nowcast quarterly GDP growth for the U.S. economy. Using the monthly FRED-MD database, we compare the nowcasting performance of five different ANN architectures: the multilayer perceptron (MLP), the one-dimensional convolutional neural network (1D CNN), the Elman recurrent neural network (RNN), the long short-term memory network (LSTM), and the gated recurrent unit (GRU). The empirical analysis presents results from two distinctively different evaluation periods. The first (2012:Q1 – 2019:Q4) is characterized by balanced economic growth, while the second (2012:Q1 – 2022:Q4) also includes periods of the COVID-19 recession. According to our results, longer input sequences result in more accurate nowcasts in periods of balanced economic growth. However, this effect ceases above a relatively low threshold value of around six quarters (eighteen months). During periods of economic turbulence (e.g., during the COVID-19 recession), longer input sequences do not help the models’ predictive performance; instead, they seem to weaken their generalization capability. Combined results from the two evaluation periods indicate that architectural features enabling long-term memory do not result in more accurate nowcasts. Comparing network architectures, the 1D CNN has proved to be a highly suitable model for GDP nowcasting. The network has shown good nowcasting performance among the competitors during the first evaluation period and achieved the overall best accuracy during the second evaluation period. Consequently, first in the literature, we propose the application of the 1D CNN for economic nowcasting.
@@
Journal of Economic Literature (JEL) codes: C45, C51, C53
@@ 1 Introduction
Gross Domestic Product (GDP) is arguably the most important flow-type monetary measure of the economic activity of a country or region. GDP is often calculated on an annual basis, but in most advanced economies it is also measured on a quarterly frequency. Considering its relevance, the availability and reliability of these quarterly data are of great significance.
@@
In our study, we apply different artificial neural networks (ANNs) to nowcast the quarterly U.S. GDP growth of the current quarter. Using the monthly FRED-MD database, we measure the nowcasting performance of five different ANN architectures relative to a naive constant growth model: The multilayer perceptron (MLP), the one-dimensional convolutional neural network (1D CNN), the Elman recurrent neural network (RNN), the long short-term memory network (LSTM), and the gated recurrent unit (GRU). These architectures are similar in that they all try to meet the above-mentioned challenges of nowcasting. However, the models have a different structure. First, the MLP’s relatively simple architecture is not explicitly designed for sequence modeling. While we can form “training sequences” by hand and add them as input, the MLP does not recognize the temporal ordering in the data. This is not to say that the MLP cannot be used effectively in applications related to time series analysis, but in this regard, the MLP is the exception. Compared to the other ANNs investigated in this paper, the MLP is a simple, time-agnostic architecture, which is not tolerant to shifts (translations) in time [37, 41]. The 1D CNN and all other RNNs investigated in this paper are specifically designed for sequence modeling. They are capable of learning order dependence in sequential prediction problems. In other words, they are trained with real input sequences. Despite this similarity, even the 1D CNN and the RNNs differ substantially in how they handle the sequential data. [^3] We will describe these differences in more detail in Section 3. From the perspective of the empirical analysis, the question comes down to what architectural features are advantageous in terms of GDP nowcasting.
@@
The first main contribution of this paper is that it proposes the utilization of a neural network architecture that has yet to be applied for economic nowcasting, namely the one-dimensional convolutional neural network (hereinafter, 1D CNN). Our results suggest that the 1D CNN could be a highly suitable architecture for GDP nowcasting as it generates accurate nowcasts for both of our evaluation periods, i.e., both during a balanced growth period and in times of high economic turbulence. While this ANN architecture has been neglected compared to the MLP and the more advanced gated RNNs, we would argue that it represents a “sweet spot” for this type of analysis in many aspects. In contrast to the MLP, the 1D CNN has an intrinsic architectural constraint that usefully limits the hypothesis space, i.e., the range of mappings the network can learn during the training process. During training, the network learns the coefficients of different finite impulse response filters, so it can only learn a composition of time-invariant (translation-invariant) mappings. Intuitively, the network performs a series of hierarchical pattern recognition tasks where the result is time-invariant to the relative position of those patterns [41]. It is worth noting that the vast majority of the traditional time series approaches (e.g., the dynamic factor model) are also built around this assumption; namely, the conditional distribution of the target variable only depends on time through the regressors (predictors). From this point of view, it seems especially advantageous to narrow the hypothesis space to this type of mappings. Unlike the MLP, the 1D CNN is trained with input sequences, reflecting the architecture’s temporal “awareness”. Training of the 1D CNN is also helped by the perception of the temporal ordering in input data, ultimately leading to less trainable parameters (weights).
@@
The major novelty of gated RNNs is that they introduce a so-called gating mechanism (practically, element-wise multiplication) that enables the network to develop long-term memory. This feature helps to overcome the vanishing gradient problem of the vanilla RNNs, making them a much more viable option for modeling time series with path-dependent behavior, i.e., with long-term memory [33, 44]. Financial time series, such as stock prices or even more foreign exchange rates, generally show characteristics of path dependency (e.g., a slowly decaying autocovariance function). Here, it can easily happen that the current value of the exchange rate is affected (supported) by a distant technical level whose effect, however, is no longer reflected in any nearby values. As a result, this old technical level should be remembered by the network in a predictive exercise. Since gated RNNs can learn those long-term dependencies between input and output, they can be highly effective in such predictive exercises. Formation of long-term memory comes at a cost, however: They contain much more trainable weights than an Elman RNN with the same number of neurons in their (hidden) state vector. Consequently, they usually need a large number of training samples to obtain a good level of generalization capability [37, 41].
@@
The second main contribution of the paper is that, first in the literature, it presents the results of a comprehensive nowcasting competition between different ANN architectures. The competitor ANN architectures are then trained with different training configurations. These training configurations differ (i) in the length of the training sequences and (ii) in information sets available intra-quarterly. The nowcasting competition presented in this paper has three important aspects and purposes. First, it aims to identify the most suitable ANN architectures for GDP nowcasting. The second research question we try to answer is how important long-term memory is in GDP nowcasting. So, we investigate how the different competitor models’ nowcasting performance changes with the length of the input sequences. Finally, we also try to measure how the nowcasting accuracy of the competitors changes with consecutive intra-quarterly data releases. So, we distinguish three different nowcasting scenarios depending on how many months of regressor data are available for the models within a given quarter. Accordingly, e.g., the 1-month nowcasting scenario refers to an information set that contains data for the monthly regressors until the first of a given quarter. We believe that the results of the empirical analysis might be indicative of some aspects of the underlying data-generating process. If, for example, the data-generating process of GDP growth was highly path-dependent (i.e., non-ergodic), then that would likely speak for the use of those gated RNNs with long input sequences.
We present the results from two evaluation (test) periods during the empirical analysis. While those evaluation periods overlap in time, they are very different in the characteristics of the target series, i.e., GDP growth. The first evaluation period ranges from 2012:Q1 to 2019:Q4, so it ends before the economic consequences of the COVID-19 pandemic hit the U.S. economy. The results for this period indicate that longer input sequences lead to more accurate nowcasts for the majority of the competitor models.
@@
However, this effect ceases above a relatively low threshold value of around six quarters (eighteen months). While the MLP works best with 8 long sequences in every nowcasting scenario, other competitors yield better accuracy with 18 long input sequences in some nowcasting scenarios. While the 1D CNN produces its best (lowest) relative RMSE and MAE values with 18 long input sequences, the overall best accuracy in the first test period is achieved by the GRU trained with 8 long sequences. Both models beat the benchmark model in terms of nowcasting performance on a 5% significance level. Our second evaluation period is from 2012:Q1 to 2022:Q4, so it also includes periods of the COVID-19 crisis. Under such circumstances, shorter input sequences lead to more accurate nowcasts for all competitor models. In the second evaluation period, the best accuracy is achieved by the 1D CNN trained with 8 long sequences. We see that architectural complexity clearly hinders the gated RNNs’ nowcasting performance in the second evaluation period. This time, they produce less accurate nowcasts than even the basic Elman RNN. Combined results from the two evaluation periods show that longer input sequences generally do not improve our models’ predictive accuracy. That is especially interesting to see in the case of the LSTM and the GRU: While their architectural design enables them to handle long input sequences, forming a long-term memory, even these networks tend to have better accuracy with shorter sequences. Considering the results from the two test periods as a whole, architectural features supporting the learning of long-term dependencies do not play an important role in GDP nowcasting. Instead, they represent a comprising trade-off between complexity and generalization capability for this specific predictive analysis.
@@ 3 Models and methods
This is not to say however, that ANNs are generally superior modeling frameworks compared to traditional linear models. While ANNs handle the curse of dimensionality more easily, they can perform worse in other areas. For example, they require a much more sophisticated data pre-processing step where we must deal with missing observations as well. One possible solution to fill up missing values is applying univariate ARMA(p,q) forecasts for each single input series. In this case, the lag lengths of the individual ARMA(p,q) models can be selected based on the value of Akaike or Schwarz information criterion [42, 51]. Although missing data issue for ANNs can be solved in many ways, state space models handle this problem by design. If we think about the three big challenges of nowcasting mentioned in the introduction, we can say that mixed-frequency data can be treated relatively well in both modeling frameworks. Compared with the DFM, proper construction of our training and test samples with mixed-frequency data may require some additional effort.
More generally, we should also mention that even the smallest ANNs have much more parameters than a typical statistical model. The large number of trainable (free) parameters results in a relatively large hypothesis space. From the bias-variance trade-off follows that with a larger hypothesis space, it is more likely to find a hypothesis which fits the training data very well. This phenomenon is called overfitting, and it happens when a model learns the fine details and noise in the training data rather than the actual systematic relationship between the regressors and the target variable(s). Overfitting can substantially weaken a statistical model’s generalization capability, and ANNs are substantially more prone to overfitting compared to the traditional time series models. So far, we tried to summarize the most important characteristics of ANNs as one of our nowcasting machines. In the following sections, we investigate those architectures in more detail that were adopted in the empirical analysis.
@@ 3.3 Recurrent neural networks
Due to vanishing gradients, RNNs tend to have a very “short” memory, limiting their usefulness in many applications [37]. Especially in those that require to learn long-term dependencies between inputs and outputs. Even with these considerations, we included one specific variant of the basic RNN in our nowcasting contest: the Elman network or Elman RNN. In an Elman network, each recurrent layer applies the following mapping for each element in the input sequence [38]:
@@
After the short description of the Elman RNN, we present how to overcome the main limitations (vanishing gradients) associated with the basic RNN by using more advanced units when building the network. Introducing our last two competitor ANN architectures, we shift our focus on more advanced RNN variants that implement a so-called gating mechanism, namely the long short-term memory network and the gated recurrent unit [33, 44].
@@
Gated recurrent unit (hereafter, GRU) represents the other gated recurrent network adopted in our nowcasting analysis. Considering its architecture, GRU is very similar to an LSTM but belongs to a newer generation of gated RNNs [33]. The GRU removes the cell state (denoted by $c_{t}$ in the LSTM) and uses its hidden state (denoted by $h_{t}$, similar to the LSTM) even for transferring information. This way, it only has two gates, a reset gate and an update gate. Thereby, GRUs has fewer parameters and tensor operations compared to LSTMs [33]. Consequently, they are a little speedier to train, and a bit less prone to overfitting. Researchers and engineers usually try both to determine which one works better for their specific use case. So we do the same in empirical analysis. Similarly to the 1D CNN, all the RNNs described above are trained with sequential data.
@@ 4 Empirical analysis
In this section, we first describe the data used in the empirical analysis and the nowcasting exercise. After that, we present the results of the empirical analysis.
@@ 4.1 Data and feature selection
The target variable of the empirical analysis is the quarterly GDP growth, that is, the percentage change in the US Gross Domestic Product relative to the preceding period. The series is measured on a quarterly frequency and is available from 1947:Q2 to 2022:Q3. Thus it contains 302 observations. Figure 6 plots the time series for quarterly GDP growth.
Figure 6: U.S. Gross Domestic Product, percent change from preceding period (quarter). Source: Federal Reserve Economic Data (FRED).
When the empirical analysis was closed, the last GDP measurement was for 2022:Q4. In our empirical analysis, we try to extract the information content of several monthly macroeconomic indicators (i.e., features) whose intra-quarterly development might be related to current economic growth. The data for the input features comes from the FRED-MD database, the monthly database for Macroeconomic Research of the Federal Reserve Bank of St. Louis, which is described extensively in [54]. [^13] The FRED-MD database is available for download under the following link: https://research.stlouisfed.org/econ/mccracken/fred-databases/.
The database consists of 134 monthly time series (indicators or features) which are clustered into eight categories:
@@
For the empirical analysis, all monthly indicators (features) are detrended using the corresponding data transformations. This is important because all of our competitor models, except for the MLP, are designed to estimate time-invariant prediction rules. Consequently, when both the target variable and the regressors (features) are stationary, we can generally expect better predictive performance from these models compared to when modeling non-stationary series. [^14] Figure A.1 below shows the level values of Real personal income (upper panel) and the result of the adequate data transformation (lower panel).
For the training of the different ANNs, we use all the monthly indicator series included in the FRED-MD database. At the same time, we introduce a bottleneck (encoder) layer, as the first layer in each network, to reduce the dimensionality of the feature space. The resulting low-dimensional embedding (state) vector contains a linear combination of the original variables, so in this respect, it performs a transformation on the data similar to principal component analysis (PCA). In the case of PCA, however, the weights of the resulting linear combination are determined to minimize the reconstruction error. By contrast, in the solution we propose, the weights of the linear combination are trained in the same integrated training process. Thus they are determined to minimize the prediction error. Considering the relatively small number of available training samples, this type of dimension reduction should significantly help each ANN’s generalization capability. Besides dimension reduction, we also apply regularization to the weight vector for the bottleneck layer ($\mathbf{w}_{b}$). Equation (6) describes the cost function augmented with the an L1 (Lasso) regularization term [47]:
@@
where $L()$ is the loss (criterion) function selected for training, $\lVert\mathbf{w}_{b}\rVert_{1}$ denotes the L1 norm of $\mathbf{w}_{b}$ and $\lambda$ is a hyperparameter which determines how severe the penalty is. [^15] We will specify the loss function in Section 4.2, during the description of the training process. Intuitively, a higher value of $\lambda$ induces a bigger penalty to the L1 norm thus it leads to a more sparse weight vector. As Equation (6) suggests, we do not use regularization in any other layer of of the different networks. This L1 (Lasso) regularization leads to a sparse weight vector in the bottleneck (encoder) layer and ultimately results in feature selection [47]. Figure 7 below shows the histograms of the estimated parameters (weights and biases) for the bottleneck layer ($\mathbf{w}_{b}^{*}$) without regularization (6(a)) and with L1 regularization(6(b)).